### v2

# Formulation of entropy through work by Carnot machine and direct derivation of law of entropy non-decrease from Kelvin principle

Abstract
We derive the law of entropy non-decrease directly from the Kelvin principle for simple and compound systems without using Clausius inequality. A key of the derivation is a formulation of entropy in terms of work by a Carnot machine operating between a system and a single heat reservoir, which is equivalent to Clausius entropy based on heat and Gyftopoulos-Beretta entropy based on work.
1 Introduction
How information of irreversibility is encoded into entropy as a state quantity? This would be one of the most difficult points for those who learn thermodynamics first to grasp. This is also true for physicists today, and the famous paper by Lieb and Yngvason was aimed to clarify the nature of entropy in terms of adiabatic accessibility in their axiomatic approach to thermodynamics [1], where neither statistical mechanics nor heat engines are used [2] (see also Ref. [3]).
Here, let us consider a conventional approach to thermodynamics. Consider a thermally isolated system in an equilibrium state $A_{1}$. After an arbitrary adiabatic process that changes the system’s state to another equilibrium state $A_{2}$, the system’s entropy $S$ never decreases before and after the change:
$S_{2}-S_{1}\geq 0,$ (1)
which is the law of entropy non-decrease, and the equality applies to reversible adiabatic processes. Here, the entropies of the system’s states $A_{1}$ and $A_{2}$ are measured by
$S_{1}=S_{0}+\int_{0}^{1}\frac{\delta Q}{T_{R}},\ \ S_{2}=S_{0}+\int_{0}^{2}\frac{\delta Q}{T_{R}},$ (2)
where $S_{0}$ is the entropy at a reference state $A_{0}$, $\delta Q$ denotes heat supplied to the system and $T_{R}$ denotes the temperature of a heat reservoir. During the process from $A_{0}$ to $A_{1}$, the heat reservoir that contacts the system and its temperature $T_{R}$ may continuously change. Equation (2) is the formula derived by Clausius. Usually, in standard textbooks on thermodynamics such as Ref. [4], after introducing Carnot cycle and Carnot efficiency as the maximum efficiency of heat engines, Eq. (1) is derived by using Clausius inequality for a cyclic process:
$\oint\frac{\delta Q}{T_{R}}\leq 0.$ (3)
The existence of entropy as a state quantity (2) is shown by using the equality condition of Clausius inequality. Then, by applying Clausius inequality Eq. (3) to a cyclic process $A_{1}\to A_{2}\stackrel{{\scriptstyle\rm rev}}{{\longrightarrow}}A_{1}$, where “rev” implies that the process is reversible, we have
$S_{2}-S_{1}-\int_{1}^{2}\frac{\delta Q}{T_{R}}\geq 0.$ (4)
If the process $A_{1}\to A_{2}$ is an adiabatic process, we may put $\delta Q=0$ into the above equality and derive Eq. (1).
The sketch of the above derivation of the law of entropy non-decrease is based on Clausius inequality Eq. (3). The Clausius inequality is derived from the Kelvin principle, which manifests one expression of the second law of thermodynamics: We cannot extract a positive work from a single heat reservoir at a fixed temperature without leaving any other changes. For a cyclic process, it can be expressed as
$W_{\rm cyc}\leq 0,$ (5)
where $W_{\rm cyc}$ denotes the work extracted from the cycle starting from an equilibrium state, and the equality applies to a reversible cycle. Deriving Clausius inequality from Kelvin principle relies on a complicated thermodynamic process tactically involving multiple heat reservoirs at different temperatures and multiple Carnot heat engines [4]. This is natural because the Kelvin principle originally refers to a process with a single heat reservoir. Moreover, the law of entropy non-decrease Eq. (1) is derived as a special case ($\delta Q=0$) of the Clausius inequality Eq. (3) and it is thus derived indirectly from the Kelvin principle.
In this paper, we show that we can derive the law of entropy non-decrease Eq. (1) directly from Kelvin principle Eq. (5) without using Clausius inequality Eq. (3). To this end, we newly define the entropy of a substance in terms of the maximum work of a Carnot machine operating between the substance and a single heat reservoir. Because we will assume only a single heat reservoir in our derivation, it apparently makes it possible to apply the Kelvin principle to show the law of entropy non-decrease. We also discuss the relationship between our entropy and Clausius entropy (2) or Gyftopoulos-Beretta entropy [5] based on work.
2 Main results
We derive the principle of entropy non-decrease Eq. (1) from Kelvin principle Eq. (5). After deriving our entropy in terms of work by a Carnot machine in Sec. 2.1, we give the derivation for simple systems in Sec 2.2 and compound systems in Sec. 2.3. Here, a simple system is a system in an equilibrium state that can be specified with the system’s temperature $T$ and volume $V$. A compound system is a system composed of simple systems.
2.1 Definition of entropy as state quantity
Figure 1: A reversible path $\gamma$ from $A_{0}$ to $A_{1}$ and a reversible path $\gamma^{\prime}$ from $A_{1}$ to $A_{0}$ on the temperature-volume ($T$$V$) plane. They consist of the adiabatic processes (arrows along the adiabatic curves) and the temperature-changing processes with a Carnot machine operating between the system at constant volume and the heat reservoir at $T_{0}$ (vertical arrows), respectively.
We introduce entropy as a state quantity in our framework. Consider a simple system initially in an equilibrium state $A_{0}$ specified with $(T_{0},V_{0})$ with $T_{0}$ and $V_{0}$ the system’s initial temperature and volume, respectively. Consider a reversible path $\gamma$ from the equilibrium state $A_{0}$ to another equilibrium state $A_{1}$ specified with $(T_{1},V_{1})$ and a reversible path $\gamma^{\prime}$ from the state $A_{1}$ to $A_{0}$ (Fig. 1). Any reversible path $\gamma$ from the equilibrium state $A_{0}$ to another one $A_{1}$ can be realized by a combination of (i) an adiabatic process (the arrow along the adiabatic curve in Fig. 1) and (ii) a temperature-changing process with a Carnot machine operating between the system at constant volume and the single heat reservoir at $T_{0}$ to raise or lower the system’s temperature (the vertical arrow in Fig. 1) by regarding the system as a finite-sized heat reservoir; any reversible path $\gamma$ may be decomposed into $N$-reversible processes $\gamma=\sum_{i=0}^{N-1}\gamma_{i}$ with $\gamma_{i}$ being the $i$-th reversible process, which is either of the above two processes (i) and (ii). In particular, when the path is a continuous one, it may be realized by a combination of the thermodynamic processes (i) and (ii) by taking the limit of $N\to\infty$.
The reversible works $W_{0\to 1}^{\rm rev}$ and $W_{1\to 0}^{\rm rev}$ extracted from $\gamma$ and $\gamma^{\prime}$, respectively, are given by
$W_{0\to 1}^{\rm rev}[\gamma]=U_{0}-U_{1}+T_{0}\int_{\gamma}\frac{\delta Q}{T},$ (6)
$W_{1\to 0}^{\rm rev}[\gamma^{\prime}]=U_{1}-U_{0}+T_{0}\int_{\gamma^{\prime}}\frac{\delta Q}{T},$ (7)
where $U$ is the internal energy and $T$ is the temperature of the system, not the temperature of the heat reservoir; the heat reservoir at fixed temperature $T_{R}=T_{0}$ is used here. We will derive Eqs. (6) and (7) later.
By applying the equality case of Kelvin principle Eq. (5) to the reversible cycle $A_{0}\stackrel{{\scriptstyle\rm rev}}{{\longrightarrow}}A_{1}\stackrel{{\scriptstyle\rm rev}}{{\longrightarrow}}A_{0}$, we have
$W_{\rm cyc}=W_{0\to 1}^{\rm rev}[\gamma]+W_{1\to 0}^{\rm rev}[\gamma^{\prime}]=0.$ (8)
From Eqs. (6)–(8), we have
$\int_{\gamma}\frac{\delta Q}{T}=-\int_{\gamma^{\prime}}\frac{\delta Q}{T}=\int_{-\gamma^{\prime}}\frac{\delta Q}{T},$ (9)
where $-\gamma^{\prime}$ denotes the reversed path of $\gamma^{\prime}$. Because $\gamma$ and $-\gamma^{\prime}$ are arbitrary paths connecting $A_{0}$ and $A_{1}$, Eq. (9) imply that the quantity $\int_{\gamma}\delta Q/T$ has a unique value independent of a path, which we denote by $\int_{0}^{1}\delta Q/T$. Therefore, we can define a state quantity $S_{1}$ at the state $A_{1}$ as
$S_{1}=S_{0}+\int_{0}^{1}\frac{\delta Q}{T}$ (10)
with $S_{0}$ at the state $A_{0}$ being a criterion value. This state quantity $S$ is our entropy. We can write the reversible work $W_{0\to 1}^{\rm rev}$ from $A_{0}$ to $A_{1}$ in Eq. (6) as
$W_{0\to 1}^{\rm rev}=U_{0}-U_{1}-T_{0}(S_{0}-S_{1})$ (11)
in a form independent of a path $\gamma$ by using the entropy. This quantity is often referred to as exergy as the maximum work extracted from a finite-sized heat reservoir until it reaches an equilibrium state with a heat reservoir at fixed temperature [6]. It should be noted that the temperature $T$ in Eq. (10) is the system’s temperature, not the heat reservoir’s one as in Eq. (2). While Eq. (2) is usually obtained via the equality case of Clausius inequality assuming thermal contact with infinitely many heat reservoirs with different temperatures, Eq. (10) has been obtained by directly applying Kelvin principle assuming only a single heat reservoir at fixed temperature $T_{0}$. But, once Eq. (10) is established, we may interpret that it is equivalent to the Clausius entropy Eq. (2), as any path in Fig. 1 can be realized with a reversible process during which a system contacts infinitely many heat reservoirs with different temperatures, and the system’s temperature $T$ agrees with the temperature $T_{R}$ of each heat reservoir for such a reversible process.
Now, we derive Eqs. (6) and (7). The work $\Delta W_{i}$ extracted from the $i$-th reversible process $\gamma_{i}$ from $(T^{(i)},V^{(i)})$ to $(T^{(i+1)},V^{(i+1)})$ as the part of $\gamma=\sum_{i=0}^{N-1}\gamma_{i}$, where $T^{(0)}=T_{0}$ and $T^{(N)}=T_{1}$, is given by the following two cases:
$({\rm i})\ \Delta W_{i}^{\rm rev}[\gamma_{i}]$ $=U(T^{(i)},V^{(i)})-U(T^{(i+1)},V^{(i+1)}),$ (12)
$({\rm ii})\ \Delta W_{i}^{\rm rev}[\gamma_{i}]$ $=U(T^{(i)},V^{(i)})-U(T^{(i+1)},V^{(i+1)})+T_{0}\int_{\gamma_{i}}\frac{\delta Q}{T},$ (13)
depending on the type of the $i$-th process, where $\Delta W_{i}^{\rm rev}[\gamma_{i}]$ in Eq. (12) for the process (i) is independent of $\gamma_{i}$ and $V^{(i+1)}=V^{(i)}$ (fixed volume) in Eq. (13) for the process (ii). These works (12) and (13) are obtained as follows: The work $\Delta W_{i}^{\rm rev}[\gamma_{i}]$ in Eq. (12) for the process (i) is just given by the internal energy change $U^{(i)}-U^{(i+1)}=U(T^{(i)},V^{(i)})-U(T^{(i+1)},V^{(i+1)})$. For the work $\Delta W_{i}^{\rm rev}[\gamma_{i}]$ in Eq. (13) for the process (ii), we operate a Carnot machine (heat engine, refrigerator, or heat pump) between the system at temperature $T$ and the heat reservoir at temperature $T_{0}$ until the system’s temperature $T$ initiating from $T^{(i)}$ reaches $T^{(i+1)}$ by keeping $V=V^{(i)}$. To raise the system’s temperature $T$, we may operate a Carnot heat engine for $T<T_{0}$ and a Carnot heat pump for $T_{0}\leq T$; to lower the system’s temperature $T$, we may operate a Carnot heat engine for $T_{0}<T$ and a Carnot refrigerator for $T\leq T_{0}$. For any case, the infinitesimal work $\delta W$ extracted by a Carnot machine operating between the system at $T$ and the heat reservoir at $T_{0}$ is commonly given by
$\delta W=-\left(1-\frac{T_{0}}{T}\right)\delta Q,$ (14)
where $\delta W$ and $\delta Q$ take a positive or negative value depending on the situation. Then, the reversible work $\Delta W_{i}^{\rm rev}[\gamma_{i}]$ in Eq. (13) extracted from the process (ii) in the total is given by the integration of Eq. (14):
$\Delta W_{i}^{\rm rev}[\gamma_{i}]=\int_{\gamma_{i}}\delta W$ $=-\int_{\gamma_{i}}\left(1-\frac{T_{0}}{T}\right)\delta Q$ (15)
$=-\int_{\gamma_{i}}\delta Q+T_{0}\int_{\gamma_{i}}\frac{\delta Q}{T}$
$=U(T^{(i)},V^{(i)})-U(T^{(i+1)},V^{(i+1)})+T_{0}\int_{\gamma_{i}}\frac{\delta Q}{T}.$
Here, in the last equality, we have used the fact that the total heat supplied to the system is equal to the internal energy change for the case of the vanishing work due to the fixed volume $V^{(i+1)}=V^{(i)}$.
By summing the work $\Delta W_{i}^{\rm rev}[\gamma_{i}]$ of each process (Eqs. (12) and (13)) along a given path $\gamma$ from $A_{0}$ to $A_{1}$, we have
$W_{0\to 1}^{\rm rev}[\gamma]$ $=\sum_{i=0}^{N-1}\Delta W_{i}^{\rm rev}[\gamma_{i}]$ (16)
$=\sum_{i=0}^{N-1}U(T^{(i)},V^{(i)})-U(T^{(i+1)},V^{(i+1)})+T_{0}\int_{\sum_{i=0}^{N-1}\gamma_{i}}\frac{\delta Q}{T}$
$=U_{0}-U_{1}+T_{0}\int_{\gamma}\frac{\delta Q}{T},$
which is Eq. (6). Equation (7) can also be calculated in the same manner.
2.2 Simple systems
Figure 2: An arbitrary adiabatic process from an equilibrium state $A_{1}$ of a simple system to another equilibrium state $A_{2}$ (dashed curve), where $A_{1}$ and $A_{2}$ are created reversibley from a reference equilibrium state $A_{0}$ (solid curves). The Kelvin principle Eq. (5) is applied to the cyclic process $A_{0}\stackrel{{\scriptstyle\rm rev}}{{\longrightarrow}}A_{1}\to A_{2}\stackrel{{\scriptstyle\rm rev}}{{\longrightarrow}}A_{0}$.
Now that we have derived the entropy in Eq. (10), we are ready to show its non-decrease during adiabatic processes. That is, we show Eq. (1) from Eq. (5). Here, we show it for simple systems.
Let a thermally isolated system be in the equilibrium state $A_{1}$ specified with $(T_{1},V_{1})$. Consider an arbitrary adiabatic process from $A_{1}$ to another equilibrium state $A_{2}$ with $(T_{2},V_{2})$ (Fig. 2) by adiabatic operations such as stirring and moving a piston attached to the system, during which the system’s state may be in nonequilibrium; after the operations, we may wait for a sufficiently long time until the system reaches the equilibrium state $A_{2}$. The work extracted from this process $W_{1\to 2}^{\rm ad}$ is given by
$W_{1\to 2}^{\rm ad}=U_{1}-U_{2},$ (17)
irrespective of whether $A_{1}\to A_{2}$ is reversible or irreversible.
Let $A_{0}$ be a reference equilibrium state of the system specified with $(T_{0},V_{0})$. We create the equilibrium state $A_{1}$ from the reference equilibrium state $A_{0}$ reversibly by combining two thermodynamic processes considered in Section 2.1 (Fig. 2). The work extracted from this process is the same as Eq. (11):
$W_{0\to 1}^{\rm rev}=U_{0}-U_{1}-T_{0}(S_{0}-S_{1}).$ (18)
Subsequently, we consider a reversible thermodynamic process from $A_{2}$ to $A_{0}$ (Fig. 2) and the reversible work $W_{2\to 0}^{\rm rev}$ extracted from it. Because this process is reversible, we may consider the process from $A_{0}$ to $A_{2}$ and reverse the sign of $W_{0\to 2}^{\rm rev}$. In the similar manner to $W_{0\to 1}^{\rm rev}$, we obtain
$W_{2\to 0}^{\rm rev}=-W_{0\to 2}^{\rm rev}=-(U_{0}-U_{2}-T_{0}(S_{0}-S_{2})).$ (19)
From Eqs. (17), (18), and (19), we obtain the work extracted from the whole cycle process $A_{0}\stackrel{{\scriptstyle\rm rev}}{{\longrightarrow}}A_{1}\to A_{2}\stackrel{{\scriptstyle\rm rev}}{{\longrightarrow}}A_{0}$ as
$W_{\rm cyc}=W_{0\to 1}^{\rm rev}+W_{1\to 2}^{\rm ad}+W_{2\to 0}^{\rm rev}=T_{0}(S_{1}-S_{2}).$ (20)
By applying Kelvin principle Eq. (5) to Eq. (20), we derive the principle of entropy non-decrease Eq. (1).
2.3 Compound systems
We show Eq. (1) from Eq. (5) for a compound system. Here, a compound system is composed of $M$ subsystems in equilibrium, where each subsystem is separated by other subsystems by insulated walls. The state of $i$-th subsystem is specified with $(T_{j},V_{j})$ ($j=1,\cdots,M$). The internal energy, entropy, and volume of the compound system are given by the sum of the ones of subsystems as $U=\sum_{j=1}^{M}U_{j}$, and $S=\sum_{j=1}^{M}S_{j}$, and $V=\sum_{j=1}^{M}V_{j}$.
In the same manner as the simple system, we consider a cycle process that consists of three thermodynamic processes $A_{1}\to A_{2}$, $A_{0}\stackrel{{\scriptstyle\rm rev}}{{\longrightarrow}}A_{1}$, and $A_{2}\stackrel{{\scriptstyle\rm rev}}{{\longrightarrow}}A_{0}$ as follows. Let $A_{1}$ be the state of a compound system specified with the states of $M_{1}$ subsystems $\{(T_{1,1},V_{1,1}),(T_{1,2},V_{1,2}),\cdots,(T_{1,M_{1}},V_{1,M_{1}})\}$ in equilibrium, where $V_{1}=\sum_{j=1}^{M_{1}}V_{1,j}$ is the volume of the compound system at $A_{1}$. Consider an arbitrary adiabatic process $A_{1}\to A_{2}$, where $A_{2}$ is specified with $M_{2}$ subsystems $\{(T_{2,1},V_{2,1}),(T_{2,2},V_{2,2}),\cdots,(T_{2,M_{2}},V_{2,M_{2}})\}$ in equilibrium. During this process, in addition to the adiabatic operations such as stirring, moving a piston attached to the system, and moving insulated walls partitioning subsystems, the number of subsystems is allowed to change from $M_{1}$ to $M_{2}$ by newly inserting or removing walls partitioning subsystems without any work cost. The work extracted from this process $W_{1\to 2}^{\rm ad}$ is given by
$W_{1\to 2}^{\rm ad}=U_{1}-U_{2},$ (21)
where $U_{1}=\sum_{j=1}^{M_{1}}U_{1,j}$ and $U_{2}=\sum_{j=1}^{M_{2}}U_{2,j}$.
Subsequently, we create the state $A_{1}$ from a reference equilibrium state $A_{0}$ of a simple system specified with $(T_{0},V_{0})$. First, we change the volume from $V_{0}$ to $V_{1}$ by a reversible adiabatic process, where we denote the state $(T_{0^{\prime}},V_{1})$ by $A_{0^{\prime}}$. The reversible work extracted from this process $A_{0}\stackrel{{\scriptstyle\rm rev}}{{\longrightarrow}}A_{0^{\prime}}$ is given by
$W_{0\to 0^{\prime}}^{\rm rev}=U_{0}-U_{0^{\prime}}.$ (22)
Then, we insert insulated walls so that the system is divided into $M_{1}$ subsystems, where each subsystem is in the equilibrium state $A_{0^{\prime},j}$ specified with $(T_{0^{\prime}},V_{1,j})$ ($j=1,\dots,M_{1}$). By operating a Carnot machine between each subsystem and the heat reservoir at temperature $T_{0}$ until it reaches the equilibrium state with $(T_{1,j},V_{1,j})$, the total reversible work extracted from this process $A_{0^{\prime}}\stackrel{{\scriptstyle\rm rev}}{{\longrightarrow}}A_{1}$ is given by
$W_{0^{\prime}\to 1}^{\rm rev}$ $=\sum_{j=1}^{M_{1}}(U_{0^{\prime},j}-U_{1,j}-T_{0}(S_{0^{\prime},j}-S_{1,j}))$ (23)
$=U_{0^{\prime}}-U_{1}-T_{0}(S_{0^{\prime}}-S_{1}).$
By summing Eqs. (22) and (23), we obtain
$W_{0\to 1}^{\rm rev}=U_{0}-U_{1}-T_{0}(S_{0^{\prime}}-S_{1}).$ (24)
Finally, the reversible work extracted from $A_{2}\stackrel{{\scriptstyle\rm rev}}{{\longrightarrow}}A_{0}$ should be calculated. Here, the system at the state $A_{2}$ consists of $M_{2}$ subsystems in equilibrium, whose state we denote by $A_{2,j}$ specified with $(T_{2,j},V_{2,j})$, where $\sum_{j=1}^{M_{2}}V_{j}=V_{2}$. In the similar manner to the calculation of $W_{0\to 1}^{\rm rev}$, we can calculate the reversible work extracted from $A_{2}\stackrel{{\scriptstyle\rm rev}}{{\longrightarrow}}A_{0}$ as the minus of $W_{0\to 2}^{\rm rev}$:
$W_{2\to 0}^{\rm rev}=-W_{0\to 2}^{\rm rev}=-(U_{0}-U_{2}-T_{0}(S_{0^{\prime\prime}}-S_{2})),$ (25)
where $S_{0^{\prime\prime}}$ is the entropy at the state $A_{0^{\prime\prime}}$ specified with $(T_{0^{\prime\prime}},V_{2})$ reached from $A_{0}$ in a reversible adiabatic process. Note that $S_{0^{\prime}}=S_{0^{\prime\prime}}$ as $A_{0^{\prime}}$ and $A_{0^{\prime\prime}}$ are on the same adiabatic curve. By summing Eqs. (21), (24), and (25), and using $S_{0^{\prime}}=S_{0^{\prime\prime}}$, we obtain the work extracted from the whole cycle process $A_{0}\stackrel{{\scriptstyle\rm rev}}{{\longrightarrow}}A_{1}\to A_{2}\stackrel{{\scriptstyle\rm rev}}{{\longrightarrow}}A_{0}$ as
$W_{\rm cyc}=W_{0\to 1}^{\rm rev}+W_{1\to 2}^{\rm ad}+W_{2\to 0}^{\rm rev}=T_{0}(S_{1}-S_{2}).$ (26)
By applying the Kelvin principle in Eq. (5) to Eq. (26), we derive the principle of entropy non-decrease in Eq. (1) for the compound system.
3 Discussion
3.1 Gyftopoulos and Beretta entropy
It should be noted that our derivation of entropy and its non-decrease in Sec. 2 shares some similarities with the approach by Gyftopoulos and Beretta (GB) [5]. By using notations in our formulation, their entropy is defined through the difference between works extracted from processes with and without the use of a heat reservoir at temperature $T_{0}$ [7]:
$S_{1}=S_{0}+\frac{W_{0\to 1}^{\rm rev}-W_{0\to 1}^{\rm ad}}{T_{0}},\ S_{2}=S_{0}+\frac{W_{0\to 2}^{\rm rev}-W_{0\to 2}^{\rm ad}}{T_{0}}.$ (27)
This definition is equivalent to Eq. (10) and is derived from it with the use of Eq. (6) by noting $W_{0\to 1}^{\rm ad}=U_{0}-U_{1}$, but the one focusing on work rather than heat. This is a natural definition as Eq. (11) suggests that the entropy serves to determine the reversible work as the maximum work extracted from a system with the aid of a heat reservoir at temperature $T_{0}$ [8]. Moreover, the GB entropy is referred to as the same entropy as the one derived by Lieb and Yngvason in their axiomatic thermodynamics [8].
By taking the difference between two in Eq. (27), we get
$S_{2}-S_{1}=\frac{W_{0\to 2}^{\rm rev}-W_{0\to 1}^{\rm rev}-W_{1\to 2}^{\rm ad}}{T_{0}},$ (28)
where $W_{1\to 2}^{\rm ad}=W_{1\to 0}^{\rm ad}+W_{0\to 2}^{\rm ad}$. Applying Kelvin principle to a cyclic process $A_{0}\stackrel{{\scriptstyle\rm rev}}{{\longrightarrow}}A_{1}\to A_{2}\stackrel{{\scriptstyle\rm rev}}{{\longrightarrow}}A_{0}$, we have
$W_{0\to 1}^{\rm rev}+W_{1\to 2}^{\rm ad}+W_{2\to 0}^{\rm rev}\leq 0\leftrightarrow-W_{0\to 1}^{\rm rev}-W_{1\to 2}^{\rm ad}+W_{0\to 2}^{\rm rev}\geq 0.$ (29)
By using Eq. (29) to Eq. (28), we obtain the law of entropy non-decrease $S_{2}-S_{1}\geq 0$. Meanwhile, it may be considered that the entropy in Eq. (27) is dependent on the choice of the reference heat reservoir through the temperature $T_{0}$. In GB approach, it is shown that Eqs. (28) and (29) are in fact independent of $T_{0}$, and the entropy is a state quantity without calculating the reversible works in Eq. (27) explicitly, and any heat reservoir at any temperature can be used for the definition of entropy as the heat reservoir merely serves as an auxiliary system (see Sec. 7.4 of Ref. [5] for their proof). In contrast, in the present approach, we evaluated the reversible works explicitly by the use of the Carnot machine and showed that the entropy is independent of the choice of the reference heat reservoir (Eq. (10)).
An advantage of GB entropy Eq. (27) to Clausius entropy Eq. (2) is that GB entropy is determined only from the measurement of $W_{0\to 1}^{\rm rev}$ for a reversible process with a heat reservoir at temperature $T_{0}$ and $W_{0\to 1}^{\rm ad}$ for an adiabatic process, with no need to measure heat. But, these two expressions are related through the expression Eq. (10).
3.2 Clausius entropy using isothermal heat
By using the Clausius entropy Eq. (2) given $S_{0}$ as the entropy of a reference state, we can calculate the entropy of another state along any reversible path. In particular, when we choose a reversible path with a single heat reservoir at temperature $T_{R}=T^{\prime}$, we have the Clausius entropy in terms of an isothermal heat:
$S_{1}=S_{0}+\frac{Q_{{\rm iso},1}}{T^{\prime}},\ \ S_{2}=S_{0}+\frac{Q_{{\rm iso},2}}{T^{\prime}},$ (30)
where $Q_{{\rm iso},1}\equiv\int_{0}^{1}\delta Q$ and $Q_{{\rm iso},2}\equiv\int_{0}^{2}\delta Q$. With this entropy in terms of the isothermal heat, it is possible to derive the law of entropy non-decrease Eq. (1) directly from the Kelvin principle by combining isothermal and adiabatic processes (see Sec. 8.11 of Ref. [9] for a similar argument): Consider that the equilibrium state $A_{1}=A_{1}(T_{1},V_{1})$ is brought to another equilibrium state $A_{2}=(T_{2},V_{2})$ adiabatically in an arbitrary manner. The equilibrium state $A_{2}$ is then brought to an equilibrium state $A_{2}^{\prime}=A_{2}^{\prime}(T^{\prime},V_{2}^{\prime})$ reversibly along the adiabatic curve. By attaching a prepared heat reservoir at temperature $T^{\prime}$ to the system, the equilibrium state $A_{2}^{\prime}$ is further brought to the equilibrium state $A_{1}^{\prime}=A_{1}^{\prime}(T^{\prime},V_{1}^{\prime})$ reversibly along the isothermal curve until it intersects the adiabatic curve with the original equilibrium state $A_{1}=A_{1}(T_{1},V_{1})$ on it. Finally, after detaching the heat reservoir from the system, the equilibrium state $A_{1}^{\prime}$ is brought to the original equilibrium state $A_{1}=A_{1}(T_{1},V_{1})$ reversibly along the adiabatic curve.
The work extracted from the whole cycle process $A_{1}\to A_{2}\stackrel{{\scriptstyle\rm rev}}{{\longrightarrow}}A_{2}^{\prime}\stackrel{{\scriptstyle\rm rev}}{{\longrightarrow}}A_{1}^{\prime}\stackrel{{\scriptstyle\rm rev}}{{\longrightarrow}}A_{1}$ is given as
$W_{\rm cyc}$ $=W_{1\to 2}^{\rm ad}+W_{2\to 2^{\prime}}^{\rm ad}+W_{2^{\prime}\to 1^{\prime}}^{\rm iso}+W_{1^{\prime}\to 1}^{\rm ad}$ (31)
$=U_{1}-U_{2}+U_{2}-U_{2}^{\prime}+W_{2^{\prime}\to 1^{\prime}}^{\rm iso}+U_{1}^{\prime}-U_{1}$
$=U_{1}^{\prime}-U_{2}^{\prime}+W_{2^{\prime}\to 1^{\prime}}^{\rm iso}$
$=U_{1}^{\prime}-U_{2}^{\prime}+U_{2}^{\prime}-U_{1}^{\prime}-T^{\prime}(S_{2}^{\prime}-S_{1}^{\prime})$
$=T^{\prime}(S_{1}^{\prime}-S_{2}^{\prime}),$
where $W_{2^{\prime}\to 1^{\prime}}^{\rm iso}=U_{2}^{\prime}-U_{1}^{\prime}-T^{\prime}(S_{2}^{\prime}-S_{1}^{\prime})$ coincides with the Helmholtz free energy difference. By applying the Kelvin principle $W_{\rm cyc}=T^{\prime}(S_{1}^{\prime}-S_{2}^{\prime})=T^{\prime}(S_{1}-S_{2})\leq 0$, we derive the law of entropy non-decrease $S_{1}\leq S_{2}$.
The outline of the above derivation is essentially the same as those in Secs. 2.2 and 2.3. An important difference is that the calculation of the work here is done by using the Helmholtz free energy difference for an isothermal process, while it is done by using the exergy for a temperature-changing process with a Carnot machine in Secs. 2.2 and 2.3. With this process using a Carnot machine, we are able to apply the Kelvin principle to the Clausius entropy Eq. (2) in terms of non-isothermal heat.
4 Conclusion
We derived the principle of entropy non-decrease Eq. (1) from Kelvin principle Eq. (5) for both simple and compound systems without using Clausius inequality Eq. (3), where the process that utilizes the Carnot machine operating between the system and the heat reservoir played the essential role in the derivation. Our entropy Eq. (10) is equivalent to Clausius entropy Eq. (2) and Gyftopoulos-Beretta entropy Eq. (27), establishing a thermodynamic consistency. We hope that the present derivation elucidates the principle of thermodynamics. It is also of interest to apply the present formulation of entropy to determine adiabatic accessibility between nonequilibrium states [8], where Gyftopoulos-Beretta entropy serves as one expression to determine such accessibility.
Acknowledgments
The author is grateful to A. Calvo Hernández for valuable comments on the earlier version of the manuscript. This work was supported by JSPS KAKENHI Grant Number 22K03450.